# Many-body critical phase in a quasiperiodic chain and dynamical Widom lines in Fock space properties

### v1

@@ I Introduction
In FS, we look into the scaling of IPR with FS dimension or the number of sites on the FS lattice $\mathcal{N}_{F}$, i.e., $IPR\simeq A_{I}\mathcal{N}_{F}^{-D_{2}}$, and the scaling of the typical value of the imaginary part of the local FS self energy, $\Delta_{t}\simeq A_{s}\mathcal{N}_{F}^{-(1-D_{s})}$. Here $D_{2}$ and $D_{s}$ are fractal and spectral fractal dimensions, respectively, and $A_{I}$ and $A_{s}$ are the amplitudes of the power laws. One expects, $D_{2}=D_{s}=1$, $D_{2}=D_{s}=0$ and $10<D_{s}<$ 1, respectively, for the delocalized, localized and non-ergodic extended states in FS. Our main results for the EAAH model, summarized in Table 1, are the following.
@@
Critical All SP excitations delocalized $A_{I}>1$, $10<D_{2}<$ 1 Anomalous $L$-dependence $\xi_{F}$ increases with $L$ (slower than MBC)
Localized All SP excitations localized $A_{I}<1$, $10<D_{2}<$ 1 $A_{s}>1$, $10<D_{s}<$ 1 (deep within the phase) $\xi_{F}$ independent of $L$ (more robust than MBL)
@@
MBC All SP excitations tend to delocalize for $L\rightarrow\infty$ $A_{I}>1$, $10<D_{2}<$ 1 Anomalous $L$-dependence $\xi_{F}$ increases with $L$
MBL All SP excitations localized $A_{I}<1$, $10<D_{2}<$ 1 $A_{s}>1$, $10<D_{s}<$ 1 $\xi_{F}$ independent of $L$
@@
(2) From the scaling of IPR, we find that critical and MBC states are indeed extended but non-ergodic, having non-zero amplitude over an infinite number, $\mathcal{N}_{F}^{D_{2}}\to\infty$ ($10<D_{2}<$ 1), albeit zero fraction of sites in the thermodynamic limit $\mathcal{N}_{F}\to\infty$. Thus, in this regard, the critical phases are similar to the multifractal states with $10<D_{2}<$ 1 in the localized and MBL phases, as well as the NEE phase Roy et al. (2023) in the GAAH model, and dissimilar to the delocalized and ergodic states, where $D_{2}=1$.
@@
(4) We find that the typical self energy $\Delta_{t}(\mathcal{N}_{F})$ does not follow any systematic scaling with $\mathcal{N}_{F}$ in the non-interacting critical and MBC phases, often exhibiting a non-monotonic dependence with $\mathcal{N}_{F}$, as well as with the parameters of the EAAH model. This is in contrast to the well-defined scaling behaviour $\Delta_{t}\sim\mathcal{N}_{F}^{-(1-D_{s})}$, seen in the delocalized and ergodic phases with $D_{s}=1$, AL and MBL phases with $10<D_{s}<$ 1, and the NEE phaseRoy et al. (2023) in the GAAH model with $10<D_{s}<$ 1.
@@
The rest of the paper is organized as follows. In Sec. II, we describe the model. In Sec. III we define the single-particle excitations in real space and then use them for characterization of the phases for noninteracting and interacting systems, in Secs. III.1 and III.2, respectively. Then in Sec. IV, we define the FS propagator and discuss the FS lattice structure. In Secs. IV.1 and IV.2, we calculate the imaginary part of the Feenberg self energy and inverse participation ratio in FS to characterize phases in noninteracting and interacting systems respectively followed by a finite size scaling of inverse participation ratio across different phase transitions in the interacting systems. Then we analyze the distributions of the FS Feenberg self-energy and inverse participation ratio in all phases in Sec. IV.3. In Sec. V, we define the FS localization length and distinguish different phases using it. Finally, we conclude in Sec. VI.
@@ II Model
Here $c_{i}$ is the fermion annihilation operator at site $i$ and $n_{i}=c^{\dagger}_{i}c_{i}$. The nearest-neighbor hopping strength $t_{i}=J+\mu\cos(2\pi\chi(i+1/2)+\phi)$ and the nearest-neighbor interaction strength $V=1$. The quasiperiodic potential $h_{i}=\lambda\cos(2\pi\chi i+\phi)$, where $\chi=(\sqrt{5}-1)/2$ and $\lambda$ is the strength of potential with a global phase $\phi\in(0,2\pi]$. The phase $\phi$ is used to generate different realizations of disorder and perform disorder averaging. By taking $\phi=0$ without loss of generality, followed by a special kind of Fourier transformation Aubry and André (1980); Biddle and Das Sarma (2010) $c_{j}=\frac{1}{\sqrt{L}}\sum_{k}e^{-i2\pi jk\chi}c_{k}$, Eq. 1 for $V=0$ in a closed chain with periodic boundary condition yields
@@
where $t^{{}^{\prime}}_{k}=\lambda/2+\mu\cos(2\pi\chi(k+1/2))$ and $h^{{}^{\prime}}_{k}=2J\cos(2\pi\chi k)$ Hence the Hamiltonians in Eq. 2 and Eq. 1 in the non-interacting limit ($V=0$) are related by a duality transformation and the model is self-dual at $\lambda=2J$ For $\mu=0$, one obtains the Aubry-André-Harper (AAH) model, with the self duality point $\lambda=2J$, which coincides with the delocalization-localization transition Aubry and André (1980). The above consideration suggests that the same AAH self-duality also exists for $\mu\neq 0$, and $\lambda=2J$ remains the delocalization to localization phase transition. This is indeed the case as shown in Fig.1(a). Moreover, the model exhibits further intriguing features for $\mu\geq J$. In this case, a distribution of zeros appears in the quasiperiodic hopping $t_{i}$’s in the thermodynamic limit such that the system can no longer support a delocalized eigenstate. This mechanism has been put forward recently as a way to produce the critical states Zhou et al. (2022). Only critical or localized eigenstates can exist in this parameter region. As a result, for small values of onsite potential strength $\lambda$ all the delocalized eigenstates turn out to be critical with $\mu=J$ being the delocalized-critical transition line for $\lambda<2J$. For $\lambda>2J$, one obtains another critical line $\lambda=2\mu$ separating the critical and localized phases, which was obtained numerically Han et al. (1994); Chang et al. (1997); Takada et al. (2004); Aditya and Roy (2024). We set $J=1$ in our calculations. The schematic of the phase diagram of the non-interacting EAAH model is shown in Fig. 1(a). The phase diagram has been obtained numerically based on earlier works Han et al. (1994); Chang et al. (1997); Takada et al. (2004); Aditya and Roy (2024) and various diagnostics used in this work, as discussed later.
@@ III Real-space single-particle excitations
Here we use local one-particle Green’s function to probe the localization properties of the single particle excitations in the real space. The single-particle Green’s function in the $n$-th many-body eigenstate |$\ket{\Psi_{n}}$ can be defined as $\mathcal{G}_{n}(i,j,t)=-i\Theta(t)\bra{\Psi_{n}}\{c_{i}(t),c^{\dagger}_{j}(0)\}\ket{\Psi_{n}}$ for sites $i$ and $j$. The Fourier transform of the local or onsite term $\mathcal{G}_{n}(i,i,t)$ can then be written as
$\mathcal{G}_{n}(i,\omega)=\sum\limits_{m}\left[\frac{|\bra{\Psi_{m}^{+}}c_{i}^{\dagger}\ket{\Psi_{n}}|^{2}}{\omega^{+}-E_{m}+E_{n}}+\frac{|\bra{\Psi_{m}^{-}}c_{i}\ket{\Psi_{n}}|^{2}}{\omega^{+}+E_{m}-E_{n}}\right],$
where |$\ket{\Psi_{m}^{+}}$ and |$\ket{\Psi_{m}^{+}}$ are the $m$-th eigenstate with energy $E_{m}$ of the system with $N+1$ and $N-1$ particles, respectively, and $\omega^{+}=\omega+\mathrm{i}\eta$, with $\omega$ the excitation energy. The broadening parameter $\eta$ is chosen to be the typical value of the many-body energy level spacing $(\sim e^{-L})$ at energy $E_{n}$ for the interacting system ($V\neq 0$) and the average single-particle energy level spacings $(\sim 1/L)$ for the noninteracting case $(V=0)$, respectively Roy et al. (2023).
In particular, we calculate the spectral function characterized by local density of states (LDOS) $\rho_{n}(i,\omega)=-(1/\pi)Im[\mathcal{G}_{n}(i,\omega)]$ for an excitation energy $\omega$ at the middle of the many-body energy spectrum ($E_{n}\simeq 0$) for the interacting EAAH model. The LDOS does not depend on many-body energy for the non-interacting case ($V=0$). We obtain the typical LDOS defined as $\ln\rho_{t}(\omega)=\langle\ln\rho_{n}(i,\omega)\rangle$ and average LDOS defined as $\rho_{a}(\omega)=\langle\rho_{n}(i,\omega)\rangle$, where delimited-⟨$\langle$..$\rangle$ … ⟩ denotes an arithmetic average over the lattice sites and $\phi$. The single particle excitations originate from the poles of the Green’s function $\mathcal{G}_{n}(i,\omega)$. For a localized state, a finite number of discrete poles, corresponding to the localization volume (length), gives rise to discrete peaks of $\mathcal{G}_{n}(i,\omega)$ in $\omega$ having zero measure in the LDOS even in the thermodynamic limit. In contrast, for a delocalized state the poles of $\mathcal{G}_{n}(i,\omega)$ form a continuum for $\omega$ lying within the energy band. Hence, $\rho_{t}(\omega)\rightarrow 0$ and $\mathcal{O}(1)$ in the thermodynamic limit for localized and delocalized phases, respectively Ganeshan et al. (2015). On the other hand, the arithmetic mean $\rho_{a}(\omega)$ approaches a non-zero value with increasing system size both in the localized and delocalized phases. Thus $\rho_{t}(\omega)$ or the ratio $\rho_{t}(\omega)/\rho_{a}(\omega)$ can be used as an order parameter to detect the localization of an excitation with energy $\omega$ Dobrosavljević et al. (2003); Dobrosavljević and Kotliar (1997); Jana et al. (2021); Roy et al. (2023).
@@ III.1 Noninteracting systems
[…]
@@
where $\psi_{k}(i)$ and $\epsilon_{k}$ are the single particle eigenfunction and energy, respectively. The $L$-dependence of both $\rho_{t}(\omega)$ and $\rho_{a}(\omega)$ can be useful to detect different types of SP excitations, delocalized, localized and gapped. Gaps are $\mathcal{O}(1)$ interval of $\omega$ in the single particle energy spectra without any spectral weight in the thermodynamic limit $L\to\infty$. Eq. 4 can be further written as $\rho(i,\omega)\approx(1/\pi\eta)\sum_{|\omega-\epsilon_{k}|\lesssim\eta}|\psi_{k}(i)|^{2}+(\eta/\pi)\sum_{|\omega-\epsilon_{k}|\gtrsim\eta}|\psi_{k}(i)|^{2}/(\omega-\epsilon_{k})^{2}$ For an excitation energy $\omega$ that falls within a gap, the second term contributes and $\rho(i,\omega)\sim|\psi_{k}(i)|^{2}/L$, where $\psi_{k}$ corresponds to state(s) at the gap edge withing the band. This leads to both $\rho_{t}(\omega)$ and $\rho_{a}(\omega)$ decreasing at least as fast as $1/L$. For excitation energy $\omega$ within an energy band, $\rho(i,\omega)\sim L|\psi_{k}(i)|^{2}+\mathcal{O}(1/L)$. Now, for delocalized and localized states $|\psi_{k}(i)|^{2}\sim 1/L$ and $|\psi_{k}(i)|^{2}\sim e^{-|i-i_{k}|/\xi}$, respectively, where $\xi$ is the single particle localization length and $i_{k}$ is localization center. As a result, for delocalized SP excitations, $\rho_{t}(\omega)$ and $\rho_{a}(\omega)$ are similar and remains nonzero in the thermodynamic limit. In contrast, the typical value of $\rho(i,\omega)$, $\rho_{t}(\omega)$, vanishes for the localized SP excitations for $L\to\infty$. However, the arithmetic mean $\rho_{a}(\omega)$ remains nonzero in the thermodynamic limit due to averaging over $i$ and/or disorder realizations, since the discrete peaks of $\rho(i,\omega)$ form a continuum when sampled over all sites and/or disorder realizations.
@@ III.2 Interacting systems
[…]
@@ IV Dynamical transitions and Widom lines in Fock-space self energy and inverse participation ratio
In this section, we study the localization properties of the different phases of the non-interacting and interacting EAAH model through the local Fock-space propagator Sutradhar et al. (2022); Roy et al. (2023), or the associated local Feenberg self energy Logan and Welsh (2019); Sutradhar et al. (2022); Roy et al. (2023), and the inverse participation ratio (IPR) of the many-body eigenstates. The Hamiltonian in Eq. 1 can be rewritten as a tight-binding model in terms of the occupation number basis {|$\{\ket{I}\}$ as Welsh and Logan (2018); Logan and Welsh (2019); Ghosh et al. (2019)
[…]
where |$\ket{I}=\ket{n_{I1}n_{I2}...n_{IL}}$ with onsite real-space occupation $n_{i}\in 0$ or $1$. The FS hopping $T_{IJ}=t_{i}$ when |$\ket{I}$ and |$\ket{J}$ are connected by a single nearest neighbor hop and $T_{IJ}=0$ otherwise. The onsite potential at the FS site $I$, given by, $E_{I}=\sum_{i}h_{i}n_{Ii}+V\sum_{i}n_{Ii}n_{Ii+1}$, generates a correlated disorder for a fictitious particle on the FS lattice or network  Welsh and Logan (2018); Logan and Welsh (2019); Ghosh et al. (2019); Roy and Logan (2020a); Altland and Micklitz (2017); Roy and Logan (2020b). The disorder-averaged distribution of the many-body density of states of the half-filled system is a Gaussian as a function of $E$, with the mean $\propto L$ and variance $\sigma_{E}^{2}=(t^{2}+\mu^{2}/2+\lambda^{2}/6+V^{2}/8)L/2$ Welsh and Logan (2018). In order to have a well-defined thermodynamic limit, we divide all the parameters by $\sqrt{L}$ and work with the rescaled hamiltonian $\tilde{H}=H/\sqrt{L}$ Welsh and Logan (2018); Roy and Logan (2020a); Logan and Welsh (2019). Since we consider here the ‘infinite-temperature state’ corresponding to the middle of the many-body energy spectrum, we set the mean energy manually to zero by implementing the transformation $\tilde{H}\rightarrow(\tilde{H}-Tr\tilde{H}/\mathcal{N}_{F})$ with $\mathcal{N}_{F}$ being the dimension of the Fock space, for each disorder realization. This is done to compensate for the fluctuations of the the middle of the spectrum with disorder realizations Sutradhar et al. (2022) for finite systems.
Figure 4: Fock-space (FS) lattice: FS lattice constructed out of real-space occupation-number basis states (black circles), illustrated for $L=8$ at half filling, starting at the top with |$\ket{11110000}$, i.e., all particles on the left side, and ending at the bottom with all particles on the right. The hoppings (black lines) and the slices (in lightblue) are indicated.
Fock-space self energy: We calculate $G_{IJ}(E)=\bra{I}G(E)\ket{J}$ for $I,J\in$ the middle (largest) slice of the FS lattice. From the diagonal elements $G_{II}(E)$, or the local FS propagator, we extract the imaginary part $\Delta_{I}(E)$ of the local self energy, or the Feenberg self energy Logan and Welsh (2019), $\Sigma_{I}(E)=X_{I}(E)-i\Delta_{I}(E)$, which is related to diagonal elements as $G_{II}(E)=[E+i\eta-E_{I}-\Sigma_{I}(E)]^{-1}$. The imaginary part of the local self-energy $\Delta_{I}(E)=Im[G_{II}^{-1}(E)]-\eta$ quantifies the inverse lifetime of a fictitious excitation created at FS site $I$ with energy $E$. Such a self energy in real space has been used earlier in the development of theory of Anderson localization Anderson (1958). We compute the typical value $\Delta_{t}$ of $\Delta_{I}(E)$, defined as $\ln\Delta_{t}=\langle\ln\Delta_{I}\rangle$, where delimited-⟨$\langle$..$\rangle$ … ⟩ represents the average over disorder realizations and FS sites $I\in$ the middle slice. In the delocalized phase $\Delta_{t}\sim\mathcal{O}(1)$, whereas in the localized or non-ergodic phases, one may expect $\Delta_{t}\rightarrow 0$ as $\mathcal{N}_{F}\rightarrow\infty$ in the thermodynamic limit. Thus $\Delta_{t}$ serves as an order parameter for the nonergodic-ergodic phase transition as found in the earlier studies Roy et al. (2023); Sutradhar et al. (2022). We expect $\Delta_{t}$ to decay as a power law, i.e., $\Delta_{t}\simeq A_{s}\mathcal{N}_{F}^{-1+D_{s}}$ with the spectral fractal dimension $D_{s}$, which is a fraction in the nonergodic phase and $1$ in the ergodic phase. As discussed later, we extract the proportionality constant $A_{s}$ and spectral dimension $D_{s}$ in the phases of the noninteracting and interacting systems. We use $16000,8000,4000,1000,500,200$ disorder realizations for system sizes $L=10,12,14,16,18,20$ respectively to generate the distributions of $\Delta_{I}$.
Fock-space inverse participation ratio: We also calculate many-body IPR in Fock-space, which has been another important quantifier for distinguishing the many-body ergodic phase from the nonergodic phases of interacting systems Macé et al. (2019); Roy et al. (2023). For a normalized eigenstate |$\ket{\Psi}=\sum_{I=1}^{\mathcal{N}_{F}}\Psi_{I}\ket{I}$, IPR is defined as $IPR=\sum_{I}|\Psi_{I}|^{4}$ $\sim A_{I}\mathcal{N}_{F}^{-D_{2}}$. The fractal dimension $D_{2}=1$ in the ergodic phase and $10<D_{2}<$ 1 in the MBL phase Macé et al. (2019). Moreover, the NEE phase in the interacting GAAH model has been reported to show $10<D_{2}<$ 1, similar to the behavior of $D_{s}$ in the same phase Roy et al. (2023). The coefficient $A_{I}$ has also been shown to be a potential indicator of the MBL transition in an earlier study Macé et al. (2019) and also related to the measure of nonergodic volume in ergodic phase Macé et al. (2019); Roy and Logan (2021). In our work, we extract the coefficient $A_{I}$ and fractal dimension $D_{2}$ from averaged IPR, calculated over disorder realizations and an energy window centered around $E=0$, for various phases in the noninteracting and interacting systems, as discussed below.
@@
Figure 6: Fock-space $IPR$ for noninteracting systems: IPR across (a) the delocalization-localization transition as a function of $\lambda$ for $\mu=0.5$, (b) delocalization-critical transition as a function of $\mu$ for $\lambda=1.0$, (c) critical-localization transitions as a function of $\lambda$ for $\mu=1.5$, and (d) localization-critical transition as a function of $\mu$ for $\lambda=3.5$. Vertical solid lines denote the points of phase transitions. In figures (c) and (d), the vertical dashed lines denote the peak/dip in IPR across the Widom-like lines in Fig.1(a) inside the critical and localized phases, respectively.

### v2

@@ I Introduction
In FS, we look into the scaling of IPR with FS dimension or the number of sites on the FS lattice $\mathcal{N}_{F}$, i.e., $IPR\simeq A_{I}\mathcal{N}_{F}^{-D_{2}}$, and the scaling of the typical value of the imaginary part of the local FS self energy, $\Delta_{t}\simeq A_{s}\mathcal{N}_{F}^{-(1-D_{s})}$. Here $D_{2}$ and $D_{s}$ are fractal and spectral fractal dimensions, respectively, and $A_{I}$ and $A_{s}$ are the amplitudes of the power laws. One expects, $D_{2}=D_{s}=1$, $D_{2}=D_{s}=0$ and $0<D_{s}<1$, respectively, for the delocalized, localized and non-ergodic extended states in FS. Our main results for the EAAH model, summarized in Table 1, are the following.
@@
Critical All SP excitations delocalized $A_{I}>1$, $0<D_{2}<1$ Anomalous $L$-dependence $\xi_{F}$ increases with $L$ (slower than MBC)
Localized All SP excitations localized $A_{I}<1$, $0<D_{2}<1$ $A_{s}>1$, $0<D_{s}<1$ (deep within the phase) $\xi_{F}$ independent of $L$ (more robust than MBL)
@@
MBC All SP excitations tend to delocalize for $L\rightarrow\infty$ $A_{I}>1$, $0<D_{2}<1$ Anomalous $L$-dependence $\xi_{F}$ increases with $L$
MBL All SP excitations localized $A_{I}<1$, $0<D_{2}<1$ $A_{s}>1$, $0<D_{s}<1$ $\xi_{F}$ independent of $L$
@@
(2) From the scaling of IPR, we find that critical and MBC states are indeed extended but non-ergodic, having non-zero amplitude over an infinite number, $\mathcal{N}_{F}^{D_{2}}\to\infty$ ($0<D_{2}<1$), albeit zero fraction of sites in the thermodynamic limit $\mathcal{N}_{F}\to\infty$. Thus, in this regard, the critical phases are similar to the multifractal states with $0<D_{2}<1$ in the localized and MBL phases, as well as the NEE phase Roy et al. (2023) in the GAAH model, and dissimilar to the delocalized and ergodic states, where $D_{2}=1$.
@@
(4) We find that the typical self energy $\Delta_{t}(\mathcal{N}_{F})$ does not follow any systematic scaling with $\mathcal{N}_{F}$ in the non-interacting critical and MBC phases, often exhibiting a non-monotonic dependence with $\mathcal{N}_{F}$, as well as with the parameters of the EAAH model. This is in contrast to the well-defined scaling behaviour $\Delta_{t}\sim\mathcal{N}_{F}^{-(1-D_{s})}$, seen in the delocalized and ergodic phases with $D_{s}=1$, AL and MBL phases with $0<D_{s}<1$, and the NEE phaseRoy et al. (2023) in the GAAH model with $0<D_{s}<1$.
@@
The rest of the paper is organized as follows. In Sec. II, we describe the model. In Sec. III we define the single-particle excitations in real space and then use them for characterization of the phases for noninteracting and interacting systems, in Secs. III.1 and III.2, respectively. Then in Sec. IV, we define the FS propagator and discuss the FS lattice structure. In Secs. IV.1 and IV.2, we calculate the imaginary part of the Feenberg self energy and inverse participation ratio in FS to characterize phases in noninteracting and interacting systems respectively followed by a finite size scaling of inverse participation ratio across different phase transitions in the interacting systems. In Sec. V, we define the FS localization length and distinguish different phases using it. Then in Sec. VI we discuss possibility of getting the Widom-like line in presence of random potential. Finally, we conclude in Sec. VII.
@@ II Model
Here $c_{i}$ is the fermion annihilation operator at site $i$ and $n_{i}=c^{\dagger}_{i}c_{i}$. We study the model at half filling. The nearest-neighbor hopping strength $t_{i}=J+\mu\cos(2\pi\chi(i+1/2)+\phi)$ and the nearest-neighbor interaction strength $V=1$. The quasiperiodic potential $h_{i}=\lambda\cos(2\pi\chi i+\phi)$, where $\chi=(\sqrt{5}-1)/2$ and $\lambda$ is the strength of potential with a global phase $\phi\in(0,2\pi]$. The phase $\phi$ is used to generate different realizations of disorder and perform disorder averaging. By taking $\phi=0$ without loss of generality, followed by a special kind of Fourier transformation Aubry and André (1980); Biddle and Das Sarma (2010) $c_{j}=\frac{1}{\sqrt{L}}\sum_{k}e^{-i2\pi jk\chi}c_{k}$, Eq. 1 for $V=0$ in a closed chain with periodic boundary condition yields
@@
where $t^{{}^{\prime}}_{k}=\lambda/2+\mu\cos(2\pi\chi(k+1/2))$ and $h^{{}^{\prime}}_{k}=2J\cos(2\pi\chi k)$ . Hence the Hamiltonians in Eq. 2 and Eq. 1 in the non-interacting limit ($V=0$) are related by a duality transformation and the model is self-dual at $\lambda=2J$ . For $\mu=0$, one obtains the Aubry-André-Harper (AAH) model, with the self duality point $\lambda=2J$, which coincides with the delocalization-localization transition Aubry and André (1980). The above consideration suggests that the same AAH self-duality also exists for $\mu\neq 0$, and $\lambda=2J$ remains the delocalization to localization phase transition. This is indeed the case as shown in Fig.1(a). Moreover, the model exhibits further intriguing features for $\mu\geq J$. In this case, a distribution of zeros appears in the quasiperiodic hopping $t_{i}$’s in the thermodynamic limit such that the system can no longer support a delocalized eigenstate. This mechanism has been put forward recently as a way to produce the critical states Zhou et al. (2022). Only critical or localized eigenstates can exist in this parameter region. As a result, for small values of onsite potential strength $\lambda$ all the delocalized eigenstates turn out to be critical with $\mu=J$ being the delocalized-critical transition line for $\lambda<2J$. For $\lambda>2J$, one obtains another critical line $\lambda=2\mu$ separating the critical and localized phases, which was obtained numerically Han et al. (1994); Chang et al. (1997); Takada et al. (2004); Aditya and Roy (2024). We set $J=1$ in our calculations. The schematic of the phase diagram of the non-interacting EAAH model is shown in Fig. 1(a). The phase diagram has been obtained numerically based on earlier works Han et al. (1994); Chang et al. (1997); Takada et al. (2004); Aditya and Roy (2024) and various diagnostics used in this work, as discussed later.
@@ III Real-space single-particle excitations
[…]
In particular, we calculate the spectral function characterized by local density of states (LDOS) $\rho_{n}(i,\omega)=-(1/\pi)Im[\mathcal{G}_{n}(i,\omega)]$ for an excitation energy $\omega$ at the middle of the many-body energy spectrum ($E_{n}\simeq 0$) for the interacting EAAH model. The LDOS does not depend on many-body energy for the non-interacting case ($V=0$). We obtain the typical LDOS defined as $\ln\rho_{t}(\omega)=\langle\ln\rho_{n}(i,\omega)\rangle$ and average LDOS defined as $\rho_{a}(\omega)=\langle\rho_{n}(i,\omega)\rangle$, where $\langle$$\rangle$ denotes an arithmetic average over the lattice sites and $\phi$. The single particle excitations originate from the poles of the Green’s function $\mathcal{G}_{n}(i,\omega)$. For a localized state, a finite number of discrete poles, corresponding to the localization volume (length), gives rise to discrete peaks of $\mathcal{G}_{n}(i,\omega)$ in $\omega$ having zero measure in the LDOS even in the thermodynamic limit. In contrast, for a delocalized state the poles of $\mathcal{G}_{n}(i,\omega)$ form a continuum for $\omega$ lying within the energy band. Hence, $\rho_{t}(\omega)\rightarrow 0$ and $\mathcal{O}(1)$ in the thermodynamic limit for localized and delocalized phases, respectively Ganeshan et al. (2015). On the other hand, the arithmetic mean $\rho_{a}(\omega)$ approaches a non-zero value with increasing system size both in the localized and delocalized phases. Thus $\rho_{t}(\omega)$ or the ratio $\rho_{t}(\omega)/\rho_{a}(\omega)$ can be used as an order parameter to detect the localization of an excitation with energy $\omega$ Dobrosavljević et al. (2003); Dobrosavljević and Kotliar (1997); Jana et al. (2021); Roy et al. (2023).
@@ III.1 Noninteracting systems
     Refer to caption(a)           Refer to caption(b)           Refer to caption(c)           Refer to caption(d)
@@
where $\psi_{k}(i)$ and $\epsilon_{k}$ are the single particle eigenfunction and energy, respectively. The $L$-dependence of both $\rho_{t}(\omega)$ and $\rho_{a}(\omega)$ can be useful to detect different types of SP excitations, delocalized, localized and gapped. Gaps are $\mathcal{O}(1)$ interval of $\omega$ in the single particle energy spectra without any spectral weight in the thermodynamic limit $L\to\infty$. Eq. 4 can be further written as $\rho(i,\omega)\approx(1/\pi\eta)\sum_{|\omega-\epsilon_{k}|\lesssim\eta}|\psi_{k}(i)|^{2}+(\eta/\pi)\sum_{|\omega-\epsilon_{k}|\gtrsim\eta}|\psi_{k}(i)|^{2}/(\omega-\epsilon_{k})^{2}$ . For an excitation energy $\omega$ that falls within a gap, the second term contributes and $\rho(i,\omega)\sim|\psi_{k}(i)|^{2}/L$, where $\psi_{k}$ corresponds to state(s) at the gap edge withing the band. This leads to both $\rho_{t}(\omega)$ and $\rho_{a}(\omega)$ decreasing at least as fast as $1/L$. For excitation energy $\omega$ within an energy band, $\rho(i,\omega)\sim L|\psi_{k}(i)|^{2}+\mathcal{O}(1/L)$. Now, for delocalized and localized states $|\psi_{k}(i)|^{2}\sim 1/L$ and $|\psi_{k}(i)|^{2}\sim e^{-|i-i_{k}|/\xi}$, respectively, where $\xi$ is the single particle localization length and $i_{k}$ is localization center. As a result, for delocalized SP excitations, $\rho_{t}(\omega)$ and $\rho_{a}(\omega)$ are similar and remains nonzero in the thermodynamic limit. In contrast, the typical value of $\rho(i,\omega)$, $\rho_{t}(\omega)$, vanishes for the localized SP excitations for $L\to\infty$. However, the arithmetic mean $\rho_{a}(\omega)$ remains nonzero in the thermodynamic limit due to averaging over $i$ and/or disorder realizations, since the discrete peaks of $\rho(i,\omega)$ form a continuum when sampled over all sites and/or disorder realizations.
@@ III.2 Interacting systems
     Refer to caption(a)           Refer to caption(b)           Refer to caption(c)           Refer to caption(d)
@@ IV Dynamical transitions and Widom lines in Fock-space self energy and inverse participation ratio
[…]
     Refer to caption(a)           Refer to caption(b)           Refer to caption(c)           Refer to caption(d)
     Refer to caption(a)           Refer to caption(b)           Refer to caption(c)           Refer to caption(d)
Figure 6: Fock-space IPRIPR for noninteracting systems: IPR across (a) the delocalization-localization transition as a function of $\lambda$ for $\mu=0.5$, (b) delocalization-critical transition as a function of $\mu$ for $\lambda=1.0$, (c) critical-localization transitions as a function of $\lambda$ for $\mu=1.5$, and (d) localization-critical transition as a function of $\mu$ for $\lambda=3.5$. Vertical solid lines denote the points of phase transitions. In figures (c) and (d), the vertical dashed lines denote the peak/dip in IPR across the Widom-like lines in Fig.1(a) inside the critical and localized phases, respectively.